For what negative value of k is there exactly one solution to this system of equations
y=2x^2+kx+6
y=-x+4

Found 2 solutions by rapture, ikleyn:
Answer by rapture(86)

Use the discriminant b^2 - 4ac.
a = 2, b = k, c = 6
Set discriminant to 0 and solve for k.
k^2 - 4(2)(6) = 0
k^2 - 8(6) = 0
k^2 - 48 = 0
k^2 = 48
sqrt{k^2} = sqrt{48}
k = sqrt{16}•sqrt{3}
k = 4•sqrt{3}, k = -4•sqrt{3}
The negative k value is -4•sqrt{3}.

Answer by ikleyn(52776)

.
The solution by the other tutor is WRONG and IRRELEVANT.

For correct solution see my post at this link
https://www.algebra.com/algebra/homework/coordinate/Linear-systems.faq.question.1085951.html

https://www.algebra.com/algebra/homework/coordinate/Linear-systems.faq.question.1085951.html